OpenStudy (anonymous):

let f(x)=x^2-7x+3 and g(x)=x+4 find f(x)+g(x)

OpenStudy (anonymous):

f(x)+g(x)=x^2-7x+3 + x+4 = x^2-6x +7

OpenStudy (anonymous):

Thank you